OpenStudy (mtalhahassan2):

Determine the velocity and acceleration of an object that moves along a straight line in such a way that its position is s(t)=t^2+(2t-3)^1/2

OpenStudy (anonymous):

Note that \[v = \frac{ ds(t) }{ dt }~~~\text{and}~~~a = \frac{ d^2s(t) }{ dt^2 } = \frac{ dv(t) }{ dt }\] so given these definitions, what would the velocity and acceleration be?

OpenStudy (mtalhahassan2):

v(t)=s(t)=2t+1/2(2t-3)^-1/2 a(t)=s1(t)=s2(t)= 2-(2t-3)-3/2

OpenStudy (anonymous):

Close, but remember you must apply the chain rule as well, \[\frac{ ds }{ dt } = 2t+\frac{ 1 }{ 2 } (2t-3)^{-1/2}*2\]

OpenStudy (mtalhahassan2):

wait where you getting that two from

OpenStudy (mtalhahassan2):

oh ok

OpenStudy (anonymous):

Everything clear/

OpenStudy (mtalhahassan2):

goted

OpenStudy (anonymous):

Alright :)

OpenStudy (mtalhahassan2):

yeah

OpenStudy (mtalhahassan2):

can you plz help me with one more question

OpenStudy (anonymous):

Sure, I'll try

OpenStudy (mtalhahassan2):

ok wait a sec

OpenStudy (mtalhahassan2):

Determine the maximum and minimum of each function on the given intervals. A) F(x)= 2x^3-9x^2, -2<x<4 B) F(x)= 12x-x^3, xzE [-3,5] C) F(x)= 2x+18/x, 1<x<5

OpenStudy (mtalhahassan2):

i wanna know how can we find the max and min in this type of situation

OpenStudy (anonymous):

First do the derivative, to find the critical points (a,b), so where would the derivative equal to 0 and is undefined, that will give you the critical points. Then you take the values and see if they are in the given interval, if they are you can continue and plug the values back into your original equation f(x) which then will give you the maximum and minimum values. Lowest value will be the minimum, highest value the maximum. I hope that was easy to follow.

OpenStudy (mtalhahassan2):

I find the 0=6x(x-3) so the critical point be 0 and 3.

OpenStudy (mtalhahassan2):

@iambatman

OpenStudy (mtalhahassan2):

@nincompoop

imqwerty (imqwerty):

yes